OpenStudy (anonymous):

Please someone help ~!!!!!!!!!!! Diana received 65 points on a project for school. She can make changes and receive three-tenths of the missing points back. She can do this 10 times . Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Diana can receive. Diana received 65 points on a project for school. She can make changes and receive three-tenths of the missing points back. She can do this 10 times . Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Diana can rece

OpenStudy (anonymous):

@Ilovecake @marylou004

OpenStudy (anonymous):

last question .

OpenStudy (ilovecake):

Hi! I am back!

OpenStudy (anonymous):

I LOVE CAKE <3 i really appreciate you

OpenStudy (ilovecake):

lol

OpenStudy (ilovecake):

use a calculator .... (3/10)^10

OpenStudy (anonymous):

0.3^10= 5.09

OpenStudy (ilovecake):

yes

OpenStudy (anonymous):

that's it lol ?

OpenStudy (ilovecake):

yup

OpenStudy (anonymous):

im going to show you what i typed

OpenStudy (ilovecake):

okay

OpenStudy (anonymous):

The formula that I have created is (3/10)^10 3/10 = 0.3 ^10 = 5.09 this shows the maximum grade Diana can receive.

OpenStudy (ilovecake):

yes. correct

OpenStudy (anonymous):

okay can you check this answer ,, then i'm done

OpenStudy (ilovecake):

alright.

OpenStudy (anonymous):

Question : Brian has been playing a game where he can create towns and help his empire expand. Each town he has allows him to create 1.17 times as many villagers. The game gave Brian 8 villagers to start with. Help Brian expand his empire by solving for how many villagers he can create with 16 towns. Then explain to Brian how to create an equation to predict the number of villagers for any number of towns. Show your work and use complete sentences.

OpenStudy (anonymous):

Answer: Ai=ai=ai= the number od villagers when there are iii towns a0-8a0=8 a0=8 villagers a1=a0*1.17 a2=a1*1.17=a0*1.17^2 ai=a0*1.17^i

OpenStudy (ilovecake):

i believe it's an exponential function. Because if you start with 8, and build 1 town, then you increase 8 by 1.17 times, but once you have a second time you increase the new amount by 1.17 times, right?

OpenStudy (ilovecake):

So therefore your answer is correct.

OpenStudy (anonymous):

THANK YOUUUUU !!!!

OpenStudy (ilovecake):

My pleasure. :)

OpenStudy (ilovecake):

What grade did you get?

OpenStudy (anonymous):

it dosen't tell me she has to grade them first , but i'll be sure to tell you!!!

OpenStudy (ilovecake):

oh. you are not in k12?